OpenStudy (anonymous):

@phi @wio @mathmale @Zale101 Which equation would best help solve the following problem? Adina has a rectangular garden that measures 9 m wide by 13 m long. She wants to increase the area to 192 m2 by increasing the width and length by the same amount. What will be the dimensions of the new garden? A. 192 = (9)(13) B. 192 = (9)(13 + x2) C. 192 = (9 – x)(13 – x) D 192 = (9 + x)(13 + x)

OpenStudy (anonymous):

D

OpenStudy (anonymous):

can you jive me the answer and also explane please

OpenStudy (anonymous):

plz explane

OpenStudy (anonymous):

D

OpenStudy (anonymous):

no medal until you explain

OpenStudy (anonymous):

and tks

OpenStudy (imstuck):

It would be D. The reason is that the formula for the area of a rectangle is A = l times w. The new area you would like to see is 192, the "A" in our formula. If you want to increase the length and the width by the same amount, keeping in mind that you don't know what the amount is, it would increase both by x. That's why.

OpenStudy (anonymous):

Area = length*breadth Let the increment in the sides be x then new length = 13+x new breadth = 9 +x new area = 192 so 192 = (13+x)*(9+x)

OpenStudy (anonymous):

i want to give one to both of you but

OpenStudy (anonymous):

thank you and i might have mor for you guys

OpenStudy (anonymous):

*more